OpenStudy (anonymous):

plz check. solve sqrt2x+15=sqrt5 x = 10 x = 5 x = –5 None of the above i got c am i correct

OpenStudy (anonymous):

please use brackets for square roots. is it sqrt(2) * x , sqrt(2x) or sqrt(2x + 15) ?

OpenStudy (anonymous):

please use proper tex

OpenStudy (anonymous):

\[\sqrt{2x+15}=\sqrt{5}\]

OpenStudy (anonymous):

this is my equation that i got the answer -5

OpenStudy (anonymous):

so basically : 2x + 15 = 5 x = -5

OpenStudy (anonymous):

yes am i correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you i just wantrd to double check myself helping my cousin with her school

OpenStudy (anonymous):

cool